OpenStudy (anonymous):

For what value of x is the rational expression below undefined? 3x-12/9-x

OpenStudy (bibby):

Any fraction/division is undefined when the denominator/divisor = 0

OpenStudy (anonymous):

so? what do i do

OpenStudy (jdoe0001):

\(\bf \cfrac{3x-12}{{\color{red}{ 9-x}}}\qquad {\color{red}{ 9-x}}=0\implies \cfrac{3x-12}{{\color{red}{ 0}}}\)

OpenStudy (jdoe0001):

so, when x = ? is the fraction undefined

OpenStudy (anonymous):

0

OpenStudy (jdoe0001):

yes, and that happens when x = ?

OpenStudy (anonymous):

so would it be 3(x-4)

OpenStudy (anonymous):

?

OpenStudy (jdoe0001):

well.. it happens if you solve for "x" at \(\bf \cfrac{3x-12}{{\color{red}{ 9-x}}}\qquad \bbox[border: 1px solid black;padding: 5px]{ {\color{red}{ 9-x}}=0}\implies \cfrac{3x-12}{{\color{red}{ 0}}}\)

OpenStudy (jdoe0001):

so.... what did you get for "x" ?

OpenStudy (jdoe0001):

9-x = 0 thus x = ?